OpenStudy (anonymous):

On a map, 3in represents 250 mi. what actual distance on this map is represented by a line 34 1/2 in long?

OpenStudy (anonymous):

divide 34 1/2 by 3 then multiply by 250

OpenStudy (anonymous):

If you want a solution to believe, \[\huge \frac 3{250} = \frac{34 \frac 12}x\] therefore \[\huge x = 34 \frac 12 \times \frac {250}3\]